# Riemannian Flow Matching for Brain Connectivity Matrices via Pullback Geometry

### v1

@@ Abstract
Generating realistic brain connectivity matrices is key to analyzing population heterogeneity in brain organization, understanding disease, and augmenting data in challenging classification problems. Functional connectivity matrices lie in constrained spaces—such as the set of symmetric positive definite or correlation matrices—that can be modeled as Riemannian manifolds. However, using Riemannian tools typically requires redefining core operations (geodesics, norms, integration), making generative modeling computationally inefficient. In this work, we propose DiffeoCFM, an approach that enables conditional flow matching (CFM) on matrix manifolds by exploiting pullback metrics induced by global diffeomorphisms on Euclidean spaces. We show that Riemannian CFM with such metrics is equivalent to applying standard CFM after data transformation. This equivalence allows efficient vector field learning, and fast sampling with standard ODE solvers. We instantiate DiffeoCFM with two different settings: the matrix logarithm for covariance matrices and the normalized Cholesky decomposition for correlation matrices. We evaluate DiffeoCFM on three large-scale fMRI datasets with more than $4600$ scans from $2800$ subjects (ADNI, ABIDE, OASIS-3) and two EEG motor imagery datasets with over $30000$ trials from $26$ subjects (BNCI2014-002 and BNCI2015-001). It enables fast training and achieves state-of-the-art performance, all while preserving manifold constraints.
@@ 1 Introduction
Modern neuroimaging analyses map brain functional signals toward connectivity matrices—covariance or correlation estimates between regions of interest or sensor channels. These structured representations are used in many applications, such as motor imagery classification [4, 43, 34, 37], brain age prediction [14, 20, 46], or disease diagnosis [13]. They are central to the analysis of signals from many neuroimaging modalities, such as functional magnetic resonance imaging (fMRI), electroencephalography (EEG), and magnetoencephalography (MEG). Brain connectivity matrices are symmetric positive definite (SPD) or lie in the open elliptope of full-rank correlation matrices, and thus belong to smooth matrix manifolds—${\mathbb{S}_{d}^{++}}$ and its submanifold ${\text{Corr}_{d}}$—defined respectively by
${\mathbb{S}_{d}^{++}}=\left\{\mathbf{\Sigma}\in\mathbb{R}^{d\times d}\,\middle|\,\mathbf{\Sigma}^{\top}=\mathbf{\Sigma},\;\mathbf{\Sigma}\succ 0\right\}\quad\text{and}\quad{\text{Corr}_{d}}=\left\{\mathbf{\Sigma}\in{\mathbb{S}_{d}^{++}}\,\middle|\,\operatorname{diag}(\mathbf{\Sigma})=\mathbf{1}\right\}.$ (1)
Several Riemannian metrics have been proposed to analyse these data such as the affine-invariant metric [54], the log-Euclidean metric [3], or the Euclidean-Cholesky metric [57]. They enable the definition of Riemannian operations such as geodesics, exponential maps, and parallel transport, which extend standard Euclidean operations to the manifold. Building on these manifolds, a wide range of machine learning algorithms have been designed for classification [4], regression [14, 20, 46], or dimension reduction [24].
@@
In this work, we address a novel and challenging issue in neuroimaging: generating realistic brain connectivity data from actual human neuroimaging data. This challenge stems from the unique structure of brain connectivity data, which is represented by SPDs or correlation matrices that lie on non-Euclidean manifolds. Moreover, neuroimaging datasets typically have limited sample sizes, making realistic data generation particularly valuable. We propose a novel Riemannian flow matching method, DiffeoCFM, based on pullback geometry, defined by a global diffeomorphism $\phi:\mathcal{M}\to E$, where $E$ is Euclidean space. This method is an efficient framework that guarantees manifold-constrained outputs by construction while avoiding computationally expensive operations specific to SPD or correlation manifolds. We instantiate DiffeoCFM with two diffeomorphisms tailored to different neuroimaging data: the matrix logarithm for SPD matrices and the normalized Cholesky decomposition for correlation matrices. Finally, we evaluate DiffeoCFM on three large-scale fMRI datasets (ADNI, ABIDE, and OASIS-3; over $4600$ scans from $2800$ subjects) and two EEG motor imagery datasets (BNCI2014-002 and BNCI2015-001; $32320$ trials from 26 subjects), demonstrating that DiffeoCFM is capable of generating realistic, neurophysiologically meaningful samples, as validated by multiple statistical metrics.
@@
[…]
@@ 2 Background
This metric induces a Riemannian norm on the tangent space $T_{x}\mathcal{M}$ at each point $x\in\mathcal{M}$: $\|\xi\|_{x}=\sqrt{(\phi^{*}g_{E})_{x}(\xi,\xi)}$. The pair $(\mathcal{M},\phi^{*}g_{E})$ is then called a pullback manifold, and $\phi^{*}g_{E}$ is the pullback metric of $g_{E}$ by $\phi$. The pullback manifold is geodesically complete and admits globally unique geodesics [57, Chap. 7]. Moreover, many Riemannian operations reduce to simple computations in $E$. Given two points $x_{0},x_{1}\in\mathcal{M}$, the geodesic $\gamma:[0,1]\to\mathcal{M}$ connecting $x_{0}$ to $x_{1}$ and its associated riemannian distance are given by
@@ 3.2 Diffeomorphic Embeddings for Generative Modeling of Brain Connectivity Matrices
To generate brain connectivity matrices, we map ${\mathbb{S}_{d}^{++}}$ and ${\text{Corr}_{d}}$ to Euclidean spaces via global diffeomorphisms: the matrix logarithm for SPD matrices and the normalized Cholesky map for correlation matrices. These maps define pullback metrics on the tangent spaces $T_{\mathbf{\Sigma}}{\mathbb{S}_{d}^{++}}={\mathbb{S}_{d}}$ and $T_{\mathbf{\Sigma}}{\text{Corr}_{d}}={\boldsymbol{\xi}\in{\mathbb{S}_{d}}\mid\operatorname{diag}(\boldsymbol{\xi})=0}$, of respective dimensions $d(d+1)/2$ and $d(d-1)/2$. This allows DiffeoCFM to perform efficient, geometry-aware generation for both matrix types.
@@ 3.3 Related work
Score-based Diffusion Models (DDPMs) [29] and CFM [39, 2] have emerged as robust, state-of-the-art baselines for generative modeling in Euclidean spaces. Several extensions have been proposed to handle data that lie on Riemannian manifolds. These include Riemannian Score-Based Generative Modeling [15], SPD-DDPM [38], and Riemannian CFM [10]. These methods tailor the loss function and ODE/SDE solvers to the geometry of a specific manifold. However, this geometric fidelity comes at a cost: manifold-specific operations such as Riemannian gradients, exponential/logarithm maps, or parallel transport must be implemented to compute the loss function and the integration on the manifold. For instance, SPD-DDPM requires a specialized neural architecture, an SPDNet, which is computationally expensive and significantly slower to train than Euclidean counterparts; see Figure 3 in Appendix for a comparison.
@@
The proposed method combines the simplicity of Euclidean training with the rigor of Riemannian geometry by using a diffeomorphism $\phi$ to embed structured matrices, apply standard CFM, and map samples back. To the best of our knowledge, it is the only method that enables generation of both SPD and correlation matrices within a unified framework. A detailed comparison of baseline methods—highlighting their assumptions, strengths, and limitations—is provided in Appendix Section G.
@@ 4.1 Metrics
To assess how well the generated data approximate the real distribution, we employ several standard quality metrics. Following Alaa et al. [1], we report $\alpha$-precision and $\beta$-recall. These metrics quantify the fidelity (how realistic the generated samples are) and the diversity (how well they span the true data distribution). We also report the harmonic mean of the two, denoted $\alpha,\beta$-F1.
@@
We follow the CAS protocol [53], training a classifier on generated samples and evaluating it on real test data. Specifically, we assess classification utility using logistic regression with cross-validated regularization, applied after mapping samples to the tangent space via the affine-invariant metric. High scores indicate that the generated data preserves task-relevant information. On fMRI datasets, the task is disease classification (control vs. patient), while for EEG, it is a two-class motor imagery problem in a brain–computer interface setting. We report ROC-AUC and F1 scores.
@@ 4.2 Datasets
The experiments include both fMRI and EEG datasets; additional details are provided in Appendix I.
@@
We use three publicly available resting-state fMRI datasets. The ABIDE dataset [45] consists of 900 subjects (one scan each), including both neurotypical and autistic individuals with a mean age of 17 years, collected across 19 international sites. The ADNI dataset [61] comprises 1,900 scans from 900 older adults (mean age 74), covering normal ageing, mild cognitive impairment, and Alzheimer’s disease. The OASIS-3 dataset [35] includes 1,000 subjects and 1,800 longitudinal sessions collected over 10 years, targeting healthy ageing and neurodegenerative conditions, with a mean participant age of 71 years. We $z$-score the time series and then compute correlation matrices (${\text{Corr}_{d}}$) using the OAS estimator [13]. We perform five random train-test splits with subject-level grouping, ensuring that scans from the same subject do not appear in both training and test sets.
@@
We use two publicly available EEG motor imagery datasets from the BCI competition. The BNCI 2014-002 dataset [56] includes 13 subjects performing right-hand and feet imagery, recorded with 15 channels over 1 session, with 80 trials per class. The BNCI 2015-001 dataset [23] comprises 12 subjects, 13 channels over 2 or 3 sessions, and 100 trials per class for the same motor imagery tasks. To process the raw data, we compute covariance matrices (${\mathbb{S}_{d}^{++}}$) using the OAS estimator [11], following standard practices [43]. In the experiments, we adopt a leave-one-session-out protocol on BNCI2014-002 and conduct cross-session experiments on BNCI2015-001.
[…]
Quality Metrics CAS Metrics $\alpha$-Precision $\uparrow$ $\beta$-Recall $\uparrow$ $\alpha$,$\beta$-F1 $\uparrow$ ROC-AUC $\uparrow$ F1 $\uparrow$ Dataset Method ABIDE Real Data 0.85 ± 0.04 0.74 ± 0.06 0.79 ± 0.08 0.67 ± 0.05 0.59 ± 0.06 DiffeoGauss 0.56 ± 0.08 0.28 ± 0.07 0.37 ± 0.11 0.67 ± 0.04 0.52 ± 0.07 TriangDDPM 0.04 ± 0.02 0.00 ± 0.00 0.00 ± 0.00 0.58 ± 0.04 0.56 ± 0.06 TriangCFM 0.06 ± 0.02 0.00 ± 0.00 0.00 ± 0.00 0.61 ± 0.04 0.42 ± 0.18 DiffeoCFM 0.74 ± 0.05 0.46 ± 0.06 0.57 ± 0.08 0.68 ± 0.03 0.60 ± 0.04 ADNI Real Data 0.88 ± 0.04 0.90 ± 0.04 0.89 ± 0.06 0.63 ± 0.04 0.64 ± 0.04 DiffeoGauss 0.03 ± 0.03 0.53 ± 0.07 0.06 ± 0.05 0.60 ± 0.06 0.16 ± 0.08 TriangDDPM 0.02 ± 0.01 0.00 ± 0.00 0.00 ± 0.01 0.54 ± 0.04 0.28 ± 0.03 TriangCFM 0.03 ± 0.00 0.00 ± 0.00 0.01 ± 0.01 0.54 ± 0.05 0.26 ± 0.15 DiffeoCFM 0.59 ± 0.15 0.77 ± 0.03 0.67 ± 0.17 0.62 ± 0.04 0.35 ± 0.09 OASIS-3 Real Data 0.86 ± 0.06 0.89 ± 0.04 0.87 ± 0.08 0.75 ± 0.04 0.64 ± 0.04 DiffeoGauss 0.49 ± 0.09 0.30 ± 0.08 0.37 ± 0.12 0.71 ± 0.07 0.45 ± 0.05 TriangDDPM 0.04 ± 0.01 0.00 ± 0.00 0.00 ± 0.00 0.52 ± 0.04 0.47 ± 0.06 TriangCFM 0.05 ± 0.01 0.00 ± 0.01 0.01 ± 0.01 0.52 ± 0.08 0.32 ± 0.19 DiffeoCFM 0.64 ± 0.08 0.38 ± 0.05 0.48 ± 0.09 0.71 ± 0.02 0.58 ± 0.03 BNCI 2014-002 Real Data 0.69 ± 0.04 0.62 ± 0.05 0.65 ± 0.07 0.83 ± 0.01 0.75 ± 0.02 DiffeoGauss 0.45 ± 0.04 0.76 ± 0.01 0.56 ± 0.05 0.80 ± 0.02 0.73 ± 0.02 TriangDDPM 0.43 ± 0.04 0.10 ± 0.02 0.16 ± 0.04 0.54 ± 0.02 0.17 ± 0.12 TriangCFM 0.48 ± 0.03 0.22 ± 0.03 0.30 ± 0.05 0.54 ± 0.03 0.21 ± 0.13 DiffeoCFM 0.63 ± 0.05 0.63 ± 0.05 0.63 ± 0.07 0.81 ± 0.02 0.73 ± 0.02 BNCI 2015-001 Real Data 0.89 ± 0.01 0.89 ± 0.01 0.89 ± 0.01 0.73 ± 0.01 0.67 ± 0.01 DiffeoGauss 0.84 ± 0.01 0.89 ± 0.01 0.87 ± 0.02 0.73 ± 0.00 0.68 ± 0.01 TriangDDPM 0.72 ± 0.03 0.55 ± 0.02 0.62 ± 0.04 0.59 ± 0.02 0.66 ± 0.02 TriangCFM 0.75 ± 0.08 0.74 ± 0.02 0.74 ± 0.08 0.62 ± 0.03 0.54 ± 0.10 DiffeoCFM 0.92 ± 0.02 0.86 ± 0.01 0.89 ± 0.02 0.73 ± 0.01 0.68 ± 0.01

### v2

@@ Abstract
Generating realistic brain connectivity matrices is key to analyzing population heterogeneity in brain organization, understanding disease, and augmenting data in challenging classification problems. Functional connectivity matrices lie in constrained spaces—such as the set of symmetric positive definite or correlation matrices—that can be modeled as Riemannian manifolds. However, using Riemannian tools typically requires redefining core operations (geodesics, norms, integration), making generative modeling computationally inefficient. In this work, we propose DiffeoCFM, an approach that enables conditional flow matching (CFM) on matrix manifolds by exploiting pullback metrics induced by global diffeomorphisms on Euclidean spaces. We show that Riemannian CFM with such metrics is equivalent to applying standard CFM after data transformation. This equivalence allows efficient vector field learning, and fast sampling with standard ODE solvers. We instantiate DiffeoCFM with two different settings: the matrix logarithm for covariance matrices and the normalized Cholesky decomposition for correlation matrices. We evaluate DiffeoCFM on three large-scale fMRI datasets with more than $4600$ scans from $2800$ subjects (ADNI, ABIDE, OASIS‑3) and two EEG motor imagery datasets with over $30000$ trials from $26$ subjects (BNCI2014‑002 and BNCI2015‑001). It enables fast training and achieves state-of-the-art performance, all while preserving manifold constraints.
Code: https://github.com/antoinecollas/DiffeoCFM
@@ 1 Introduction
[…]
${\mathbb{S}_{d}^{++}}=\left\{\mathbf{\Sigma}\in\mathbb{R}^{d\times d}\,\middle|\,\mathbf{\Sigma}^{\top}=\mathbf{\Sigma},\mathbf{\Sigma}\succ 0\right\}\text{and }{\text{Corr}_{d}}=\left\{\mathbf{\Sigma}\in{\mathbb{S}_{d}^{++}}\,\middle|\,\operatorname{diag}(\mathbf{\Sigma})=\mathbf{1}\right\}.$ (1)
Several Riemannian metrics have been proposed to analyse these data such as the affine-invariant metric [65, 7], the log-Euclidean metric [3], or the Euclidean-Cholesky metric [68]. They enable the definition of Riemannian operations such as geodesics, exponential maps, and parallel transport, which extend standard Euclidean operations to the manifold. Building on these manifolds, a wide range of machine learning algorithms have been designed for classification [4], regression [16, 22, 55], or dimension reduction [27].
@@
In this work, we address a novel and challenging issue in neuroimaging: generating realistic brain connectivity data from actual human neuroimaging data. This challenge stems from the unique structure of brain connectivity data, which is represented by SPDs or correlation matrices that lie on non-Euclidean manifolds. Moreover, neuroimaging datasets typically have limited sample sizes, making realistic data generation particularly valuable. We propose a novel Riemannian flow matching method, DiffeoCFM, based on pullback geometry, defined by a global diffeomorphism $\phi:\mathcal{M}\to E$, where $E$ is Euclidean space. This method is an efficient framework that guarantees manifold-constrained outputs by construction while avoiding computationally expensive operations specific to SPD or correlation manifolds. We instantiate DiffeoCFM with two diffeomorphisms tailored to different neuroimaging data: the matrix logarithm for SPD matrices and the normalized Cholesky decomposition for correlation matrices. Finally, we evaluate DiffeoCFM on three large-scale fMRI datasets (ADNI, ABIDE, and OASIS-3; over $4600$ scans from $2800$ subjects) and two EEG motor imagery datasets (BNCI2014-002 and BNCI2015-001; $30000$ trials from $26$ subjects), demonstrating that DiffeoCFM is capable of generating realistic, neurophysiologically meaningful samples, as validated by multiple statistical metrics.
@@
$\mathcal{M}$ is a smooth manifold with tangent space $T_{x}\mathcal{M}$ and Riemannian norm $\|\cdot\|_{x}$. We write $\gamma$ for geodesics, and $\dot{\gamma}(t)\triangleq\frac{d}{dt}\gamma(t)$ for curve’s speed. Let $\phi:\mathcal{M}\to E$ be a global diffeomorphism to Euclidean space $E$, with differential $\operatorname{D}\phi(x):T_{x}\mathcal{M}\mapsto E$ and inverse $(\operatorname{D}\phi(x))^{-1}$. The pushforward of a distribution $p$ on $\mathcal{M}$ is $\phi_{\#}p$, defined via $\int f\,d(\phi_{\#}p)=\int(f\circ\phi)\,dp$ for any $f$ continuous on $E$. We denote $x\mid y\sim p(\cdot\mid y)$ for conditional sampling with label $y\in\mathcal{Y}$. Let $\mathbb{S}_{d}$} be the space of symmetric $d\times d$ matrices, $\mathbb{S}_{d}^{++}$} the SPD cone, and $\text{Corr}_{d}$} the set of correlation matrices. Let $\text{LT}_{d}^{1}$ be the set of lower-triangular matrices with unit diagonal. We define $\mathrm{vec}_{\mathrm{sl}}:{\mathbb{S}_{d}}\mapsto\mathbb{R}^{d(d-1)/2}$ as the operator extracting strictly lower-triangular entries, and $\mathrm{vec}_{\mathrm{lt}}$ for the full lower-triangular part (diagonal included), with $\sqrt{2}$ scaling off-diagonal terms.
@@ 2 Background
This metric induces a Riemannian norm on the tangent space $T_{x}\mathcal{M}$ at each point $x\in\mathcal{M}$: $\|\xi\|_{x}=\sqrt{(\phi^{*}g_{E})_{x}(\xi,\xi)}$. The pair $(\mathcal{M},\phi^{*}g_{E})$ is then called a pullback manifold, and $\phi^{*}g_{E}$ is the pullback metric of $g_{E}$ by $\phi$. The pullback manifold is geodesically complete and admits globally unique geodesics [68, Chap. 7]. Moreover, many Riemannian operations reduce to simple computations in $E$. Given two points $x_{0},x_{1}\in\mathcal{M}$, the geodesic $\gamma:[0,1]\to\mathcal{M}$ connecting $x_{0}$ to $x_{1}$ and its associated riemannian distance are given by
@@ 3.2 Diffeomorphic Embeddings for Generative Modeling of Brain Connectivity Matrices
To generate brain connectivity matrices, we map $\mathbb{S}_{d}^{++}$} and $\text{Corr}_{d}$} to Euclidean spaces via global diffeomorphisms: the matrix logarithm for SPD matrices and the normalized Cholesky map for correlation matrices. These maps define pullback metrics on the tangent spaces $T_{\mathbf{\Sigma}}{\mathbb{S}_{d}^{++}}={\mathbb{S}_{d}}$ and $T_{\mathbf{\Sigma}}{\text{Corr}_{d}}=\{\boldsymbol{\xi}\in{\mathbb{S}_{d}}\mid\operatorname{diag}(\boldsymbol{\xi})=0\}$, of respective dimensions $d(d+1)/2$ and d(d1)/2d(d-1)/2. This allows DiffeoCFM to perform efficient, geometry-aware generation for both matrix types. Note that we selected these two diffeomorphisms for their simplicity and ease of implementation. However, other choices are possible. For correlation matrices, alternative parameterizations are discussed in [68, Chapter 7]. For SPD matrices, one can use the Cholesky factor with a logarithm applied to the diagonal, leading to the log-Cholesky metric [46], which also defines a global diffeomorphism.
@@ 3.3 Related work
Denoising Diffusion Probabilistic Models (DDPMs) [31] and CFM [47, 2] have emerged as robust, state-of-the-art generative models in Euclidean spaces. Several extensions have been proposed to handle data that lie on Riemannian manifolds. These include Riemannian Score-Based Generative Modeling [17], SPD-DDPM [45], and Riemannian CFM [12]. These methods tailor the loss function and ODE/SDE solvers to the geometry of a specific manifold. However, this geometric fidelity comes at a cost: manifold-specific operations such as Riemannian gradients, exponential/logarithm maps, or parallel transport must be implemented to compute the loss function and the integration on the manifold. For instance, SPD-DDPM requires a specialized neural architecture, an SPDNet, which is computationally expensive and significantly slower to train than Euclidean counterparts; see Figure 4 in Appendix for a comparison. References [35] and [38] explore more general settings by learning bridge matches on arbitrary manifolds or data-driven Riemannian metrics, whereas our approach focuses on Riemannian geometries defined via known pullback diffeomorphisms. In [26], the authors leverage reparameterisation with normalising flows to learn probability densities on Lie groups (non-Euclidean spaces), and thus can be seen as an early precursor to our approach.
@@
The proposed method combines the simplicity of Euclidean training with the rigor of Riemannian geometry by using a diffeomorphism $\phi$ to embed structured matrices, apply standard CFM, and map samples back. To the best of our knowledge, it is the only method that enables generation of both SPD and correlation matrices within a unified framework. A detailed comparison of baseline methods—highlighting their assumptions, strengths, and limitations—is provided in Appendix Section H.
@@ 4.1 Metrics
The metrics used are method-agnostic; that is, they are computed solely from the generated samples. They fall into two categories: (i) quality metrics, which assess how well the generated data approximate the real data distribution; and (ii) classification accuracy score (CAS) metrics, which evaluate the usefulness of generated data for training classifiers.
@@
We assess how closely the generated samples align with the real data distribution using the $\alpha$-precision and $\beta$-recall metrics introduced by [1]. We compute these metrics using the EvaGeM library [^1] https://github.com/nicolassalvy/EvaGeM. In contrast to [1], which uses a Deep neural network, EvaGeM employs a One-Class SVM, providing a more stable and hyperparameter-robust estimator across datasets. These metrics quantify the fidelity (how realistic the generated samples are) and the diversity (how well they span the true data distribution). We also report the harmonic mean of the two, denoted $\alpha,\beta$-F1.
@@
We follow the CAS protocol [64], training a classifier on generated samples and evaluating it on real test data. Specifically, we assess classification utility using a logistic regression with liblinear solver, balances class weights and a 5-fold cross-validation to select the inverse regularization strength $C$ from the grid $\{10^{-4},10^{-3},\dotsc,10^{4}\}$. High scores indicate that the generated data preserves task-relevant information. On fMRI datasets, the task is disease classification (control vs. patient), while for EEG, it is a two-class motor imagery problem in a brain–computer interface setting. We report ROC-AUC and F1 scores.
@@ 4.2 Datasets
The experiments include both fMRI and EEG datasets; additional details are provided in Appendix G.
@@
We use three publicly available resting-state fMRI datasets. The ABIDE dataset [53] consists of 900 subjects (one scan each), including both neurotypical and autistic individuals with a mean age of 17 years, collected across 19 international sites. The ADNI dataset [73] comprises 1,900 scans from 900 older adults (mean age 74), covering normal ageing, mild cognitive impairment, and Alzheimer’s disease. The OASIS-3 dataset [42] includes 1,000 subjects and 1,800 longitudinal sessions collected over 10 years, targeting healthy ageing and neurodegenerative conditions, with a mean participant age of 71 years. We $z$-score the time series and then compute correlation matrices ($\text{Corr}_{d}$}) using the OAS estimator [15]. We report mean and standard deviations computed across 10 random train-test splits with subject-level grouping, ensuring that scans from the same subject do not appear in both training and test sets.
@@
We use two publicly available EEG motor imagery datasets from the BCI competition. The BNCI2014-002 dataset [67] includes $13$ subjects performing right-hand and feet imagery, recorded with 15 channels over $1$ session, with $80$ trials per class. The BNCI2015-001 dataset [25] comprises $12$ subjects, $13$ channels over $2$ or $3$ sessions, and $100$ trials per class for the same motor imagery tasks. To process the raw data, we compute covariance matrices ($\mathbb{S}_{d}^{++}$}) using the OAS estimator [13], following standard practices [51]. We report performance on a leave-one-session-out protocol on BNCI2014-002 and on cross-session experiments on BNCI2015-001. The reported standard deviations are computed over 5 inner splits and averaged over sessions.